OpenStudy (anonymous):

If the (a+1) th , 7th , and (b+1) th terms of an A.P are in G.P with a ,6 , b being in H.P then 4th term of this A.p is?

OpenStudy (anonymous):

Can you rewrite this without abbreviations?

OpenStudy (anonymous):

it should be there to denote the term "th"

OpenStudy (anonymous):

lol...i cant understand what this question says !!

hartnn (hartnn):

i am getting answer as 0 !!....do u have options??? @Yahoo!

OpenStudy (anonymous):

The answer should be 0

OpenStudy (anonymous):

Yup u r correct

OpenStudy (anonymous):

Can u show that @hartnn

hartnn (hartnn):

Quite simple if u have tried: 3(a+b)=ab......HP condition let A be 1st term.,D be common difference 'n'th term is A+(n-1)D so,GP condition gives (2A+6D)^2=(2A+aD)(2A+bD) 4A^2 +36D^2+24AD=4A^2+abD^2+2AD(a+b) on simplifying this, (36-ab)D=2A(a+b-12) then use HP condition so that 36-ab cancels out on both sides and what remains is -3D=2A--->4th term=0

OpenStudy (anonymous):

i tried lol!

hartnn (hartnn):

now u got it? @Yahoo!

OpenStudy (anonymous):

,GP condition gives (2A+6D)^2=(2A+aD)(2A+bD)???????

OpenStudy (anonymous):

b^2 = ac

OpenStudy (anonymous):

t7 = a + 6d

OpenStudy (anonymous):

u have taken it as 2A

OpenStudy (anonymous):

@hartnn r u there

hartnn (hartnn):

oh my mistake,i am reconsidering now,and will reply u very soon

OpenStudy (anonymous):

ok

hartnn (hartnn):

so,GP condition gives (A+6D)^2=(A+aD)(A+bD) A^2 +36D^2+12AD=A^2+abD^2+AD(a+b) on simplifying this, (36-ab)D=A(a+b-12) then use HP condition so that 36-ab cancels out on both sides and what remains is -3D=A--->4th term=0 sorry for previous mistake,still i get 4th term=0 :P

OpenStudy (anonymous):

HP condition = 3(a+b)=ab

OpenStudy (anonymous):

hw did u use that

OpenStudy (anonymous):

ok...i got it

OpenStudy (anonymous):

thxxx @hartnn